OpenStudy (anonymous):

Can someone help me find the solution to this equation 4t(t-5)-2t(3t+2)+108=3t(2t-5)-t(8t-3)

OpenStudy (anonymous):

4t(t-5)-2t(3t+2)+108=3t(2t-5)-t(8t-3) distribute 4t^2-20t-6t-4t+108= 6t^2-15t-8t^2+3t combine like terms -2t^2-24t+108=-2t^2-12 try to do the rest youself... honestly, I got stuck figuring it out from here. Sorry :(

OpenStudy (anonymous):

Thank you so much! I love you leandra